find three consecutive odd integers such that twice the sum of the first two 11 is more than three times the greatest

Answer by CubeyThePenguin(3113)

consecutive odd integers: (x-2), x, (x+2)

2((x-2) + x) = 11 + 3(x+2)
2(2x-2) = 3x + 17
4x - 4 = 3x + 17
x = 21

The integers are 19, 21, and 23.